Circular No. 5898

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

(2060) CHIRON M. W. Buie, Lowell Observatory, reports: “The stellar occultation

candidate Ch02, identified by Bus et al. (1993, submitted to A.J.), was predicted by C. Olkin, S. McDonald, C. Ford, J. Foust, L. Sopata, and J. L. Elliot with sufficient accuracy that observations by four teams were successful in observing the occultation by (2060) Chiron on Nov. 7. These teams were manned by the University of Arizona, Massachusetts Institute of Technology, Ames Research Center, NASA, and Lowell Observatory at the following locations: Tierra Del Sol, CA (R. Marcialis and R. Hill); Palomar Mountain (S. Bus and R. Meserole); Ojai, CA (L. Wasserman and J. Spencer); and Table Mountain (E. Dunham, M. Buie, and J. Young). J-band images were also obtained around the time of the event by T. H. Jarrett, C. A. Beichman, and T. Herter using the Palomar Prime Focus Infrared Camera on the 5-m Hale Telescope. The flux of the star was seen to drop by 80 percent for 7 s at the Tierra Del Sol site. The observation is consistent with a single chord, placing a lower bound of 166 km for the diameter of Chiron. However, this interpretation requires that the star be an unresolved double of which only one component was occulted by Chiron. Preliminary spectra from the Perkins 1.8-m telescope do not confirm the duplicity of the star, but the results are not yet conclusive. Each of the other sites noted shallower unresolved dips that decrease in strength with increasing distance from Chiron and occur at nearly the same time, but they are inconsistent with chords from a solid surface occultation. These dips may be due to larger-than-average noise, or to an occultation by a narrow (< 10 km) dust jet.”

C. Telesco, Marshall Space Flight Center, NASA; B. Schulz, ESTEC, European Space Agency; and H. Campins and D. Osip, University of Florida, communicate: “We report 10- and 20-μm detections of (2060) Chiron obtained on Nov. 22 UT at the NASA Infrared Telescope Facility with the MSFC 20-pixel bolometer array (pixel size: α × δ ∼ 4′′.3 × 4′′.2) during exceptionally good atmospheric conditions. The following flux densities do not include color corrections due to the low temperature of Chiron: Nov. 22.6217-22.6388, 20 ± 5 mJy at 10.8 μm (Δλ = 5.3 μm; this is the first detection of Chiron at this wavelength); Nov. 22.6425-22.6492, 310 ± 61 mJy at 19.2 μm (Δλ = 5.2 μm). Uncertainties in absolute calibration, which are not included above, should not exceed 10 percent.”

1993 November 29 Daniel W. E. Green